# Incommensurate and commensurate antiferromagnetic orders in the kagome compound UV6Sn6

### v1

@@ IV Conclusion
Our study reveals that the UV6Sn6 structure deviates from the HfFe6Ge6-type structure commonly reported for $RT_{6}X_{6}$ systems. The structure is best described as a $7\times 7$ superstructure resulting from a commensurate modulation of the HfFe6Ge6-type parent cell on the $ab$ plane, driven by an ordering pattern of the uranium atoms. We propose a model that accurately replicates the observed features in X-ray and neutron diffraction data.
Magnetic characterization from magnetization and neutron diffraction measurements demonstrate two successive magnetic transitions at $T_{N1}$ = 30 K and $T_{N2}$ = 23.5 K. The first transition corresponds to an incommensurate antiferromagnetic ordering that locks to a commensurate antiferromagnetic ground state at $T_{N2}$. This behavior suggests competing exchange interactions. The slightly large effective moment ($3.89(1)$$\mu_{B}$/U) and strong uniaxial anisotropy along the c-axis point to significant spin-orbit coupling, a hallmark of uranium-based systems.
Our work invites future investigations concerning the relationship between the newly discovered UV6Sn6 superstructure and the electronic band structure (e.g. via ARPES or DFT calculations). Additionally, further investigations such as pressure or doping effects might shed light on the origin of the structural and magnetic complexity in this new member of the 166 family.
@@ IV.1 Acknowledgements
We thank Eric Bauer, Priscila Rosa, Sean Thomas for letting us know about their upcoming Arxiv posting on the same compound. M.A.P. thanks F. Denis Romero for fruitful discussions about crystallography. M.A.P. and V.T. acknowledge funding from the Laboratoire d’excellence LANEF in Grenoble (ANR-10-LABX-51-01) and the support from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of the Office of International Science and Engineering (OISE). P.L.B. was supported by QuantEdu-France (ANR-22-CMAS-0001). We further acknowledge financial support from the French National Agency for Research ANR within the Project FRESCO (ANR-20-CE30-0020) and SCATE (ANR-22-CE30-0040).
The neutron scattering data collected at ILL for the present work are available at Ref. 44 for D23 and Ref. 45 for XtremeD.
@@ Appendix A Appendix A: Details of structural refinement of UV6Sn6 using a HfFe6Ge6-type model
Appendix B Appendix B: The goodness of fit and weighed profile R-factor
The goodness-of-fit factor, GOF, is given by [46]:
$\text{GOF}=\sqrt{\frac{\sum_{i}\left(\frac{y_{i}^{\text{obs}}-y_{i}^{\text{calc}}}{\sigma_{i}}\right)^{2}}{\nu}}$
Where:
  • $y_{i}^{\text{obs}}$ is the observed intensity at the $i$-th data point.
  • $y_{i}^{\text{calc}}$ is the calculated intensity at the $i$-th data point from the model.
  • $\sigma_{i}$ is the standard deviation or uncertainty of the observed intensity.
  • $\nu$ are the degrees of freedom during the refinement, given by:
    $\nu=N-P$
    Where $N$ is the total number of data points, and $P$ is the number of refined parameters.
This parameter compares the deviations between observed and calculated intensities relative to the expected noise and it is normalized by the degrees of freedom allowed during the refinement. The best possible value for GOF is 1, with values close to unity indicating that the differences between the observed and calculated data are consistent with the experimental noise.
The weighed profile $R$-factor, $wR$, can be written as [46]:
$wR=\sqrt{\frac{\sum_{i}w_{i}\left(y_{i}^{\text{obs}}-y_{i}^{\text{calc}}\right)^{2}}{\sum_{i}w_{i}\left(y_{i}^{\text{obs}}\right)^{2}}}$
With weight:
$w_{i}=\frac{1}{\sigma_{i}}$
The parameter $wR$ is not normalized by the total uncertainty throughout the data set. This leads to $wR$ being more sensitive than GOF to deviations between observed and calculated intensities. In particular, for weak reflections, such as satellites, the observed intensity $y_{i}^{\text{obs}}$ is small, but the weight $w_{i}$ becomes large if the experimental $\sigma_{i}$ is small (as is our case with the single-crystal XRD experiments). This means that any small mismatch between $y_{i}^{\text{obs}}$ and $y_{i}^{\text{calc}}$ will contribute significantly to the numerator of $wR$, inflating its value.
Appendix C Appendix C: Evidence of superstructure in neutron diffraction
As mentioned in the main text, the superstructure peaks associated with the UV6Sn6 supercell were first observed using XtremD. Fig. 9 shows representative data for the satellites associated with the (2, 0, -3) reflection (in the HfFe6Ge6-type cell) obtained from the XtremD detector bank at 27 K. These satellites were also observed on D23 with the same positions at 24 K and 150 K, which rules out their link to the magnetic ordering below 30 K evidenced in this work.
@@ Appendix C Appendix C: Evidence of superstructure in neutron diffraction
Appendix D Appendix D: Refinement using a YCo6Ge6-type model
The YCo6Ge6-type structure is related to the HfFe6Ge6-type by a $c/2$ translation, and a halving of the unit cell along the $c$-axis and of the occupancy at the U and Sn atomic positions, as illustrated in Fig. 10 (positions $1a$ and $2c$ in Table 3). For this model, the HfFe6Ge6-type-cell odd-$l$ planes are no longer integers. This means that Bragg reflections are not expected at those values of $l$ during refinement. Although refinement of this YCo6Ge6-type model against X-ray diffraction data leads to a reduced $wR$ (Table 3), it does not accurately describe the UV6Sn6 structure as discussed in the main text.
@@ Appendix D Appendix D: Refinement using a YCo6Ge6-type model
Appendix E Appendix E: $RT_{6}X_{6}$ supercells
@@
Figure 11: The CoSn-type structure is formed by alternating layers of (a) Sn honeycomb lattices and (b) Sn-centered Co kagome networks (Sn in gray and Co in green). The resulting structure depicted in (c) –note two unit cells are stacked along $c$, leads to the HfFe6Ge6-type structure when a rare-earth is introduced in (a), at the origin of the unit cell (marked in blue). The surrounding atoms at this position form an hexagonal cavity represented as a polyhedron in (c). Electrostatic and steric (size-related) effects, impose preferences regarding the simultaneous occupation of neighboring cavities along $c$ [22]. These preferences may lead to patterns that extend the unit cell into a supercell such as (d).

### v3

@@ IV Conclusion
Our study reveals that the UV6Sn6 structure deviates from the HfFe6Ge6-type structure commonly reported for $RT_{6}X_{6}$ systems. The structure is best described as a $7\times 7\times 1$ modulation of the HfFe6Ge6-type parent cell on the $ab$ plane, driven by an ordering pattern of the uranium atoms. This model replicates the observed features in x-ray and neutron diffraction data.
Magnetic characterization from magnetization and neutron diffraction measurements demonstrate two successive magnetic transitions at $T_{N1}\approx 29$ K and $T_{N2}$ = 23.5 K. The first transition corresponds to an incommensurate antiferromagnetic ordering that locks to a commensurate antiferromagnetic ground state at $T_{N2}$. This behavior suggests competing exchange interactions. The slightly large effective moment ($3.89(1)$$\mu_{B}$/U) and strong uniaxial anisotropy along the $c$ axis point to significant spin-orbit coupling, a hallmark of uranium-based systems.
Our paper opens several future research avenues. The superstructure of UV6Sn6 emphasizes the unique role of uranium’s $5f$ orbitals. The complexity of the crystallographic and magnetic structures invites further investigation. Exploring pressure or chemical doping could help clarify the interplay between uranium’s $5f$-driven structural modulations and the vanadium kagome network in UV6Sn6. Applying pressure may suppress the $7\times 7\times 1$ supercell, providing insights into the relationship between steric and electronic contributions to the modulation, and shed some light on competing orders. Doping may tune the hybridization between the $5f$ and $3d$ orbitals in the kagome network. Such studies would advance our understanding of both the structural complexity of UV6Sn6 and the broader class of 166 kagome compounds.
@@ IV.1 Acknowledgements
We thank E. Bauer, P. Rosa, S. Thomas for letting us know about Ref. [33] as well as for sharing additional x-ray diffraction data. M.A.P. thanks F. Denis Romero for fruitful discussions about crystallography. M.A.P. and V.T. acknowledge funding from the Laboratoire d’excellence LANEF in Grenoble (ANR-10-LABX-51-01) and the support from the US National Science Foundation (NSF) Grant No. 2201516 under the Accelnet program of the Office of International Science and Engineering (OISE). P.L.B. was supported by QuantEdu-France (ANR-22-CMAS-0001). We further acknowledge financial support from the French National Agency for Research ANR within Project FRESCO (ANR-20-CE30-0020) and SCATE (ANR-22-CE30-0040).
IV.2 Data availability
The data that support the findings of this article are openly available [53, 54]. The rest of the data are available upon reasonable request from the authors.
@@ Appendix A Appendix A: Details of structural refinement of UV6Sn6 using a HfFe6Ge6-type model
Appendix B Appendix B: Evidence of superstructure in neutron diffraction
As mentioned in the main text, the superstructure peaks associated with the UV6Sn6 supercell were first observed using XtremD. Figure 9 shows representative data for the satellites associated with the (2, 0, -3) reflection (in the HfFe6Ge6-type cell) obtained from the XtremD detector bank at 27 K. These satellites were also observed on D23 with the same positions at 24 K and 150 K, which rules out their link to the magnetic ordering below 29 K evidenced in this paper.
@@ Appendix B Appendix B: Evidence of superstructure in neutron diffraction
Appendix C Appendix C: Refinement using a YCo6Ge6-type model
The YCo6Ge6-type structure is related to the HfFe6Ge6-type by a $c/2$ translation, and a halving of the unit cell along the $c$ axis and of the occupancy at the U and Sn atomic positions, as illustrated in Fig. 10 (positions $1a$ and $2c$ in Table 3). For this model, the HfFe6Ge6-type cell odd-$l$ planes are no longer integers. This means that Bragg reflections are not expected at those values of $l$ during refinement. Although refinement of this YCo6Ge6-type model against x-ray diffraction data leads to a reduced $wR2$ (Table 3), it does not accurately describe the UV6Sn6 structure as discussed in the main text.
@@ Appendix C Appendix C: Refinement using a YCo6Ge6-type model
Appendix D Appendix D: $RT_{6}X_{6}$ supercells
@@
Figure 11: The CoSn-type structure is formed by alternating layers of (a) Sn honeycomb lattices and (b) Sn-centered Co kagome networks (Sn in gray and Co in green). The resulting structure depicted in (c)–note two unit cells are stacked along $c$–leads to the HfFe6Ge6-type structure when a rare-earth is introduced in (a), at the origin of the unit cell (marked in blue). The surrounding atoms at this position form a hexagonal cavity represented as a polyhedron in (c). Electrostatic and steric (size-related) effects impose preferences regarding the simultaneous occupation of neighboring cavities along $c$ [22]. These preferences may lead to patterns that extend the unit cell into a supercell such as (d).
Appendix E Appendix E: Comparison of $T_{N}$
Figure 12: Determination of transition temperatures $T_{N1}$ and $T_{N2}$ by the temperature derivative of the magnetization for $H\parallel c$ (red) measured at 1 T, the resistivity (green, 0 T), and specific heat (blue, 0 T).
Figure 12, presents a comparison of the temperature dependence of the magnetization derivative, $dM/dT$ –for a magnetic field applied along the $c$ axis, the resistivity derivative, $d\rho/dT$, and the specific heat divided by temperature, $C_{p}/T$. The arrows mark the magnetic transition temperatures $T_{N1}$ and $T_{N2}$. Rather good agreement is observed between the different measurements. The only slight discrepancy is that $T_{N2}$, as determined from the resistivity, appears slightly higher. This difference may arise from the use of a different single-crystal sample for that particular measurement.